### 21 Mar 2026 v1

# Redshift Dipoles from Non-Geodesic Observer Congruences in Covariant Cosmology

Abstract
In an inhomogeneous universe, the Hubble frame used to describe the cosmic expansion does not, in general, coincide with a geodesic matter flow. In this work, it is shown that within a fully covariant framework, a non-geodesic observer congruence introduces an additional contribution to the relation between affine parameter and observed redshift, proportional to the line-of-sight projection of the observer four-acceleration. This induces a dipolar modulation of the redshift itself, which propagates to any observable expressed in redshift space.
Unlike the standard kinematic dipole associated with a global Lorentz boost, this contribution is sourced by the kinematics of the observer congruence along the past light cone and can exhibit a non-trivial redshift dependence. These results provide a direct framework to identify such signatures in cosmological data, with potential implications for the interpretation of large-scale dipoles and redshift-based inference.
1 Introduction
Cosmological observables are defined on the past light–cone of the observer, reflecting the causal structure of spacetime [5, 6]. Among them, redshift plays a central role, encoding the relation between spacetime dynamics and observational data. In standard cosmology, it is interpreted with respect to the comoving geodesic congruence, for which it reduces to the familiar relation $1+z=a_{0}/a$.
However, this identification implicitly assumes a particular choice of observer congruence. In realistic, inhomogeneous spacetimes, the congruence used to define cosmological observables—such as those associated with the Hubble flow or the CMB rest frame—need not coincide with the geodesic matter frame. In general, different congruences are related by spacetime-dependent local boosts, under which geodesicity is not preserved. As a result, an observer congruence may possess a non-vanishing four-acceleration even in a spacetime sourced by pressureless matter. This raises a fundamental question: how does the choice of observer congruence affect the definition and interpretation of cosmological observables?
In this work, the problem is addressed within a fully covariant framework [5, 12]. The optical focusing equation governing the angular diameter distance is purely geometrical, while the relation between affine parameter and observed redshift depends explicitly on the kinematics of the observer congruence. In particular, a non-geodesic congruence introduces additional contributions to the redshift propagation equation, leading to a dipolar modulation proportional to the projection of the four-acceleration along the line of sight. This result has immediate observational consequences. Since cosmological distances are expressed as functions of redshift, any dipolar modulation in the redshift propagates to all redshift-space observables. Importantly, this dipole originates from the local kinematics of the observer congruence and is not, in general, reducible to a single global Lorentz boost.
An interesting observational point is that accelerated reference frames are not merely a theoretical possibility. The acceleration of the Solar System barycentre produces the well-known secular aberration drift [11, 2], a dipolar proper-motion pattern that has been measured with very long baseline interferometry and more recently constrained with Gaia astrometry. This shows that observer acceleration can leave directly observable anisotropic signatures in precision cosmological measurements, and it motivates considering whether analogous kinematic effects may also enter the definition of redshift when the relevant cosmological frame is not geodesic. More broadly, several large-scale dipolar or directional anomalies have been discussed in the literature, including the low-multipole alignments in the CMB and the unexpectedly large dipole reported in quasar number counts (see e.g. [7, 10, 3]). While such observations need not be explained by the mechanism proposed here, they illustrate that the interpretation of cosmic dipoles remains open and deserves careful scrutiny. In that context, the possibility that part of the observed anisotropy may arise from the kinematics of a non-geodesic observer congruence is worth considering, precisely because it does not rely on new degrees of freedom or modifications of gravity, but follows from the covariant relativistic description itself.
The structure of this paper is as follows. The Sachs optical equation for the angular diameter distance is first derived, and its purely geometrical character is emphasized. The propagation of redshift in both geodesic and non-geodesic observer congruences is then analyzed, showing that a dipolar modulation naturally arises in the latter case. Order-of-magnitude estimates for the expected amplitude of this effect are presented, and an operational observable to isolate the dipolar component in redshift data is introduced. The paper concludes with a discussion of the implications for cosmological inference and possible observational tests.
2 The angular diameter distance from the Raychaudhuri equation
Consider a congruence of null geodesics with tangent vector $k^{\alpha}$, affinely parametrised by $\lambda$,
$k^{\beta}\nabla_{\beta}k^{\alpha}=0,\qquad k^{\alpha}k_{\alpha}=0.$ (2.1)
Let $A(\lambda)$ denote the cross-sectional area of a narrow bundle of rays, defined on the two-dimensional screen space orthogonal to $k^{\alpha}$. The associated null expansion scalar is
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (2.2)
We define the optical shear and vorticity scalars of the null congruence as
$\sigma_{N}^{2}\equiv\frac{1}{2}\,\sigma^{(N)}_{\alpha\beta}\sigma_{(N)}^{\alpha\beta}\geq 0,\qquad\omega_{N}^{2}\equiv\frac{1}{2}\,\omega^{(N)}_{\alpha\beta}\omega_{(N)}^{\alpha\beta}\geq 0,$ (2.3)
so that $\sigma^{(N)}_{\alpha\beta}\sigma_{(N)}^{\alpha\beta}=2\sigma_{N}^{2}$ and $\omega^{(N)}_{\alpha\beta}\omega_{(N)}^{\alpha\beta}=2\omega_{N}^{2}$.
The evolution of the null expansion $\theta_{N}$ is governed by the null Raychaudhuri equation [8],
$\frac{d\theta_{N}}{d\lambda}=-\frac{1}{2}\theta_{N}^{2}-2\sigma_{N}^{2}+2\omega_{N}^{2}-R_{\alpha\beta}k^{\alpha}k^{\beta}.$ (2.4)
With this convention, $\sigma_{N}^{2}$ contributes to focusing, while $\omega_{N}^{2}$ contributes to defocusing.
For a hypersurface-orthogonal null congruence (as appropriate for light bundles forming wavefronts), the vorticity vanishes,
$\omega_{N}^{2}=0.$ (2.5)
Using $\theta_{N}=(1/A)\,dA/d\lambda$ and the identity (see Appendix A)
$\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\theta_{N}^{2}=\frac{2}{\sqrt{A}}\frac{d^{2}\sqrt{A}}{d\lambda^{2}},$
Raychaudhuri can be rewritten as an evolution equation for $\sqrt{A}$:
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}\left(R_{\alpha\beta}k^{\alpha}k^{\beta}+2\sigma_{N}^{2}\right)\sqrt{A}.$ (2.6)
For a shear-free light bundle,
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\sqrt{A}.$ (2.7)
The angular diameter distance is defined geometrically by
$d_{A}^{2}\equiv\frac{A}{\Omega},$ (2.8)
where $\Omega$ is the solid angle subtended by the source at the observer. The solid angle is defined in the observer’s local rest frame at the vertex of the light cone and is therefore independent of the affine parameter along the null geodesic. Hence $d_{A}\propto\sqrt{A}$, and the evolution equation for $\sqrt{A}$ in the shear-free case reduces to
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A}.$ (2.9)
Equation (2.9) corresponds to the shear-free limit of the Sachs optical equation [9], valid for hypersurface-orthogonal null congruences with vanishing vorticity. In this case, the evolution of the angular diameter distance is governed solely by the Ricci focusing term. The initial conditions at the observer (located at $\lambda=0$) are
$d_{A}(0)=0,\qquad\left.\frac{dd_{A}}{d\lambda}\right|_{0}=1,$ (2.10)
where the second condition fixes the normalization of the affine parameter such that, locally near the vertex of the light cone, $d_{A}\simeq\lambda$ and $A\simeq\lambda^{2}$. This reflects the locally flat behaviour of the light cone near the observer.
2.1 Matter source term and FLRW universe
Using Einstein’s equations,
$R_{\alpha\beta}=8\pi G\left(T_{\alpha\beta}-\frac{1}{2}Tg_{\alpha\beta}\right)+\Lambda g_{\alpha\beta},$ (2.11)
and noting that $g_{\alpha\beta}k^{\alpha}k^{\beta}=0$, we obtain
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G\,T_{\alpha\beta}k^{\alpha}k^{\beta}.$ (2.12)
For a perfect fluid,
$T_{\alpha\beta}=(\rho+p)u_{\alpha}u_{\beta}+pg_{\alpha\beta}.$ (2.13)
Here $u^{\alpha}$ denotes the four-velocity of the cosmological perfect fluid sourcing the geometry. This implies
$T_{\alpha\beta}k^{\alpha}k^{\beta}=(\rho+p)(u\cdot k)^{2}.$ (2.14)
Thus the focusing term becomes
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)(u\cdot k)^{2}.$ (2.15)
Note that $(u\cdot k)$ is the energy $E$ of the photon measured by the observer $u^{\alpha}$ and therefore
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)E^{2}.$ (2.16)
As a consistency check, specialising to an FLRW spacetime where the null congruence associated with radial light propagation is shear-free and the metric takes its standard form, the solution of Eq. (2.9) is (see Appendix B)
$d_{A}=a(\eta)\,S_{K}\!\big(\chi\big).$ (2.17)
with $\eta$ being the conformal time and $\chi$ the comoving distance.
3 Redshift definition and the geodesic observer congruence
From this point on, $u^{\alpha}$ will denote the observer congruence used to define measured photon energy and redshift; in exact FLRW it coincides with the fluid 4-velocity. The observed redshift is defined covariantly by
$1+z_{obs}\equiv\frac{(u\cdot k)_{\rm em}}{(u\cdot k)_{\rm obs}},\qquad E\equiv-(u\cdot k),$ (3.1)
Differentiating $E=-(u\cdot k)$ along the ray gives
$\frac{dE}{d\lambda}=-\,k^{\alpha}\nabla_{\alpha}(u_{\beta}k^{\beta})=-\,k^{\alpha}k^{\beta}\nabla_{\alpha}u_{\beta},$ (3.2)
where we used $k^{\alpha}\nabla_{\alpha}k^{\beta}=0$ (affine null geodesic). Introducing the standard 1+3 decomposition of $\nabla_{\alpha}u_{\beta}$:
$\nabla_{\alpha}u_{\beta}=\frac{1}{3}\theta\,h_{\alpha\beta}+\sigma_{\alpha\beta}+\omega_{\alpha\beta}-u_{\alpha}\mathcal{A}_{\beta},\qquad h_{\alpha\beta}=g_{\alpha\beta}+u_{\alpha}u_{\beta}.$ (3.3)
For the photon, decompose $k^{\alpha}$ relative to the observers as
$k^{\alpha}=E\,(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,$ (3.4)
where $e^{\alpha}$ is the spatial propagation direction measured by $u^{\alpha}$.
Substituting (3.3) and (3.4) into (3.2), using $k^{\alpha}k_{\alpha}=0$ and $h_{\alpha\beta}u^{\beta}=0$, one obtains the exact kinematical propagation law:
$\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta+\sigma_{\alpha\beta}e^{\alpha}e^{\beta}-\mathcal{A}_{\alpha}e^{\alpha}\right).$ (3.5)
For a geodesic observer congruence,
$\mathcal{A}^{\alpha}\equiv u^{\beta}\nabla_{\beta}u^{\alpha}=0.$ (3.6)
In the absence of shear in the observer congruence (as in exact FLRW), we have
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,\frac{1}{3}E\theta.}$ (3.7)
Equation (3.7) reduces to the standard FLRW redshift law. Indeed, defining the expansion scalar and local Hubble rate of the observer congruence by
$H\equiv\frac{\theta}{3},$ (3.8)
one has
$\frac{1}{E}\frac{dE}{d\lambda}=-EH.$ (3.9)
We define a redshift parameter along the null geodesic as $1+z(\lambda)=E(\lambda)/E_{\rm obs}$ along the ray whose value at the source position $\lambda=\lambda_{s}$ coincides with the observed redshift $z_{obs}$ and $E_{\rm obs}$ is constant at the observation event. It follows that
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH.$ (3.10)
On the other hand, in FLRW the scale factor depends only on cosmic time (see Appendix C), so
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha},\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (3.11)
Therefore,
$\frac{da}{d\lambda}=k^{\alpha}\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}k^{\alpha}=\dot{a}\,E=aHE.$ (3.12)
where we used $u_{\alpha}k^{\alpha}=-E$ and $H=\dot{a}/a$. Then
$\frac{1}{a}\frac{da}{d\lambda}=EH.$ (3.13)
Comparing (3.10) and (3.13), we obtain
$\frac{d\ln(1+z)}{d\lambda}=-\frac{d\ln a}{d\lambda},$ (3.14)
Evaluating at the observer, the last equation integrates to
$1+z_{obs}=\frac{a_{0}}{a},$ (3.15)
where $a$ is the scale factor evaluated at the source position. Thus, Eq. (3.5) provides the exact covariant propagation law whose geodesic, shear-free FLRW limit reproduces the standard cosmological redshift relation.
4 Non-geodesic observer congruence and the dipole modulation
The Sachs equation (2.9) for $d_{A}$ is purely geometric and does not require the observer congruence to be geodesic. In the FLRW derivation above, the assumption of geodesicity entered only when we related the affine parameter (or conformal time) to the redshift parameter $z$ defined along the null geodesic. We now relax this assumption and make explicit where and how the modification appears.
In the previous FLRW derivation we implicitly took $u^{\alpha}$ to be the comoving geodesic congruence. We now relax (3.6) and allow
$\mathcal{A}^{\alpha}\neq 0.$ (4.1)
The impact of (4.1) is entirely contained in the propagation of $E$ along the null ray. Neglecting the shear of the observer congruence for simplicity, the propagation law reduces to
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta-\mathcal{A}_{\alpha}e^{\alpha}\right).}$ (4.2)
Note that the $\mathcal{A}_{\alpha}e^{\alpha}$ term is direction-dependent. Equation (4.2) can be written as
$\frac{1}{E}\frac{dE}{d\lambda}=-EH+E\,\mathcal{A}_{\alpha}e^{\alpha},$ (4.3)
where $H=\theta/3$. Using $1+z=E/E_{\rm obs}$, with $E_{\rm obs}$ constant, we obtain
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH+E\,\mathcal{A}_{\alpha}e^{\alpha}.$ (4.4)
At this stage we keep the FLRW background relation for the scale factor evolution 3.13, so that the modification arises entirely from the observer-congruence dependence of the redshift propagation law. We obtain
$\frac{d}{d\lambda}\ln\!\big[(1+z)a\big]=E\,\mathcal{A}_{\alpha}e^{\alpha}.$ (4.5)
Therefore, the standard relation $(1+z)a=a_{0}$ is replaced by
$(1+z_{obs})=\frac{a_{0}}{a}\exp\left(\int_{\lambda_{0}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right).$ (4.6)
For weak non-geodesic effects, this becomes
$1+z_{obs}\simeq\frac{a_{0}}{a}\left[1+\int_{\lambda_{0}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right].$ (4.7)
Since $\mathcal{A}_{\alpha}e^{\alpha}$ is the projection of a spatial vector along the line of sight, it has the angular structure of a dipole. This shows that non-geodesicity of the observer congruence naturally induces a dipolar modulation in the redshift itself. The non-geodesic dipole arises from a line-of-sight integral of the observer acceleration field and therefore depends on the photon trajectory and redshift. This contrasts with the standard kinematic dipole, which is generated by a constant boost and is independent of distance. As a result, this dipole cannot, in general, be removed by a single global Lorentz transformation, as it arises from an integrated, position-dependent acceleration field rather than a constant boost.
5 Order–of–Magnitude Estimate of the Non–Geodesic Dipole
The dipolar contribution associated with a non–geodesic observer congruence arises from the term $\mathcal{A}_{a}e^{a}$ in the covariant propagation equation of the redshift,
$\frac{d}{d\lambda}\ln(1+z)=-E\left(\frac{1}{3}\theta-\mathcal{A}_{a}e^{a}\right).$ (5.1)
A simple estimate can be obtained by isolating the non-geodesic contribution. We focus on the dipolar term alone and define
$\mathcal{A}_{\parallel}\equiv\mathcal{A}_{a}e^{a}.$ (5.2)
The shear term has already been neglected in the previous section, so that only the isotropic expansion term and the dipolar non–geodesic term remain. To estimate the amplitude of the anisotropic modulation, it is convenient to separate the isotropic monopole contribution as
$\ln(1+z)\simeq\ln(1+\bar{z})+\delta\ln(1+z)_{\rm ng}.$ (5.3)
where $\bar{z}$ is the associated monopolar redshift, whose evolution is
$\frac{d}{d\lambda}\ln(1+\bar{z})=-EH.$ (5.4)
The dipolar residual is therefore governed by
$\left.\frac{d}{d\lambda}\,\delta\ln(1+z)\right|_{\rm ng}\simeq E\,\mathcal{A}_{\parallel},\qquad\mathcal{A}_{\parallel}\equiv\mathcal{A}_{a}e^{a}.$ (5.5)
To express this in a more intuitive form, recall that along the photon trajectory
$k^{a}=\frac{dx^{a}}{d\lambda},$ (5.6)
and therefore, in the local rest frame of the observer congruence,
$E=-u_{a}k^{a}=\frac{dt}{d\lambda},$ (5.7)
up to the usual choice of units and sign convention. [^1] More precisely, in the local rest frame of the observer congruence one has $E=c\,dt/d\lambda$ if $c$ is kept explicit. Using
$\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}=E\frac{d}{dt},$ (5.8)
Eq. (5.5) becomes
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\simeq\mathcal{A}_{\parallel},$ (5.9)
in units with $c=1$. Restoring dimensions, the relevant rate is $\mathcal{A}_{\parallel}/c$, so that
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{c}.$ (5.10)
Integrating along the photon path gives
$\delta\ln(1+z_{obs})_{\rm ng}\sim\int\frac{\mathcal{A}_{\parallel}}{c}\,dt.$ (5.11)
If $\mathcal{A}_{\parallel}$ varies slowly along the line of sight over the redshift range of interest, this reduces to
$\delta\ln(1+z_{obs})_{\rm ng}\approx\frac{\mathcal{A}_{\parallel}}{c}\,\Delta t\approx\frac{\mathcal{A}_{\parallel}L}{c^{2}},$ (5.12)
where $L\simeq c\,\Delta t$ is the effective light–travel distance.
At low redshift, the light-travel distance is of order the Hubble-law distance,
$L\simeq\frac{c}{H_{0}}\,z.$ (5.13)
Substituting Eq. (5.13) into Eq. (5.12) yields
$\delta\ln(1+z_{obs})_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (5.14)
For $z\ll 1$, one has $\delta\ln(1+z)\simeq\delta z$ to leading order, so the same estimate can be written directly as
$\delta z_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (5.15)
Inverting this relation gives an estimate of the effective acceleration required to generate a dipole of amplitude $\delta z_{\rm ng}$ at redshift $z$:
$\mathcal{A}_{\parallel}\sim\frac{cH_{0}}{z}\,\delta z_{\rm ng}.$ (5.16)
Taking $H_{0}\simeq 70\,{\rm km\,s^{-1}\,Mpc^{-1}}$, one finds
$cH_{0}\simeq 6.8\times 10^{-10}\ {\rm m\,s^{-2}}.$ (5.17)
At $z\simeq 0.1$, representative of the upper end of the shallow redshift regime relevant for CosmicFlows analyses (see [14]), Eq. (5.16) becomes
$\mathcal{A}_{\parallel}\simeq 6.8\times 10^{-9}\,\delta z_{\rm ng}\quad{\rm m\,s^{-2}}.$ (5.18)
Therefore, a dipole amplitude of $\delta z_{\rm ng}\sim 10^{-3}$, $\delta z_{\rm ng}\sim 10^{-4}$ and $\delta z_{\rm ng}\sim 10^{-5}$ requires, respectively
$\mathcal{A}_{\parallel}\sim 7\times 10^{-12}\ {\rm m\,s^{-2}},$ (5.19)
$\mathcal{A}_{\parallel}\sim 7\times 10^{-13}\ {\rm m\,s^{-2}},$ (5.20)
$\mathcal{A}_{\parallel}\sim 7\times 10^{-14}\ {\rm m\,s^{-2}}.$ (5.21)
For comparison, a characteristic cosmological acceleration associated with the buildup of a peculiar velocity $v\sim 300\,{\rm km\,s^{-1}}$ over a Hubble time is
$\mathcal{A}_{\rm H}\sim vH_{0}\sim 7\times 10^{-13}\ {\rm m\,s^{-2}},$ (5.22)
which lies in the range required to generate a dipole of order $\delta z\sim 10^{-4}$ at $z\sim 0.1$. Here $\mathcal{A}_{\rm H}$ should be understood as an order-of-magnitude estimate for the spatial component of the observer 4–acceleration, expressed in terms of the typical rate of change of peculiar velocities over a Hubble time.
This suggests that the required scale is not obviously negligible from a cosmological point of view, but rather comparable to cosmological acceleration scales associated with structure formation. The main discriminant is therefore not the amplitude alone, but the redshift evolution and directional stability of the dipole.
6 Observational Implementation and Testable Signatures
In a perturbed FLRW spacetime, peculiar velocities of sources and observers produce a Doppler dipole. If a source moves with a small peculiar velocity $v_{s}^{\alpha}$ relative to the cosmological frame $u^{\alpha}$, and the observer has velocity $v_{o}^{\alpha}$, the observed redshift at linear order in peculiar velocities can be written as
$1+z\simeq(1+\bar{z})\left[1+(v_{s}-v_{o})_{\alpha}e^{\alpha}\right],$ (6.1)
where $\bar{z}$ denotes the purely monopolar redshift generated by the background expansion.
Combining this with the covariant propagation law derived above and for weak non-geodesic effects, the dipolar component of the redshift field can schematically be written as
$\delta z\sim(1+\bar{z})\left[(v_{s}-v_{o})_{\alpha}e^{\alpha}+\int_{\lambda_{o}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right],$ (6.2)
where the first term represents the standard kinematic contribution from peculiar velocities, while the second term arises from the non–geodesic nature of the observer congruence.
Both contributions share the same angular structure (dipolar), but differ in their physical origin and redshift dependence. The kinematic term is local and reflects peculiar velocities at emission and observation, whereas the non–geodesic contribution is an integrated effect along the photon trajectory. This distinction provides the basis for an observational test. In peculiar velocity surveys such as CosmicFlows [14], redshifts are typically corrected to the CMB rest frame, removing the dominant contribution from the observer’s local motion. In this case, the remaining dipole is expected to be dominated by large–scale structure (LSS) kinematics, with a possible additional contribution arising from non–geodesic effects.
At low redshift, this motivates a phenomenological model for the dipolar modulation of the redshift field of the form
$\delta z(\bar{z})=\,\hat{n}\cdot\mathbf{D}(\bar{z}),$ (6.3)
where $\bar{z}$ is the isotropic (monopole) component and $\mathbf{D}(\bar{z})=D(\bar{z})\,\hat{p}(\bar{z})$ is the dipole vector, with $D(\bar{z})$ its amplitude and $\hat{p}(\bar{z})$ its direction.
In this framework, the standard kinematic dipole is expected to reflect the coherence scale and amplitude of peculiar velocities, and therefore to exhibit a redshift dependence determined by large–scale structure dynamics. In contrast, a non–geodesic contribution introduces an additional component sourced by a line–of–sight integral of the observer acceleration field, which need not follow the same scaling. As a consequence, the two contributions need not follow the same redshift evolution. Phenomenologically, one may write
$\mathbf{D}(\bar{z})=\mathbf{D}_{\rm kin}(\bar{z})+\mathbf{D}_{\rm ng}(\bar{z}),$ (6.4)
where the two contributions need not be aligned and need not follow the same redshift evolution.
Therefore, by measuring the redshift dependence of the dipole amplitude and its directional stability, one can test whether the observed signal is consistent with purely kinematic expectations or requires an additional non–geodesic contribution. This provides a concrete observational avenue to test whether the redshift dipole is fully accounted for by large–scale structure kinematics, or whether additional contributions are present.
7 Discussion and Conclusions
In this work, it has been shown that non-geodesic observer congruences generically induce a dipolar modulation in the cosmological redshift through the projection of the four-acceleration along the line of sight. This effect arises at the level of the redshift itself and therefore propagates to any observable expressed in redshift space. The observable is defined operationally through redshift measurements in a specified frame (e.g. CMB frame), and therefore the effect corresponds to a physically distinct inference, rather than a coordinate artifact or a trivial reparametrization.
A key aspect of this result is that the induced dipole is not equivalent to the standard kinematic dipole generated by a constant peculiar velocity. In the latter case, the dipole can be removed by a single global Lorentz boost. In contrast, the contribution derived here is sourced by the (generally position-dependent) acceleration of the observer congruence and can accumulate along the past lightcone, making it generically non-reducible to a global boost transformation and leading to a genuinely non-local contribution to the observed redshift.
This point is closely related, but not equivalent, to previous work on tilted cosmologies, where observer congruences moving with respect to the matter flow exhibit modified kinematical properties (see e.g. [13, 1]). In such approaches, the focus is typically on local frame transformations and their impact on dynamical variables. By contrast, the effect derived here arises from the propagation of redshift along the past lightcone and therefore encodes an integrated contribution of the observer acceleration, which is not captured by local boosts alone.
This distinction has direct implications for cosmological inference. In practice, observational data are interpreted in terms of redshifts defined with respect to specific frames, such as the CMB frame. If such frames correspond to non-geodesic congruences, the inferred cosmological parameters may inherit systematic anisotropies or biases associated with observer kinematics. Standard treatments of redshift corrections in peculiar-velocity analyses (e.g. [4]) typically assume geodesic observer congruences, and therefore do not account for additional line-of-sight contributions associated with non-geodesic effects.
The order-of-magnitude estimates presented here indicate that the acceleration required to produce observable dipolar effects is comparable to typical cosmological acceleration scales associated with large-scale structure formation. This suggests that the effect may be relevant in low-redshift datasets, where peculiar velocities and bulk flows are significant. In this regime, the observable of interest is not only the dipole amplitude itself, but its redshift evolution and directional stability, which can in principle distinguish between kinematic and non-geodesic contributions. In particular, large-scale dipoles inferred from redshift surveys (e.g. [15]) provide a natural observational benchmark against which such effects could be tested. While these measurements are commonly interpreted in terms of bulk flows induced by large-scale structure, the framework presented here suggests that part of the observed signal could, in principle, receive contributions from the kinematics of the observer congruence itself.
A quantitative assessment requires extending the present framework to realistic cosmological models and datasets, including forward modeling of redshift-space observables and their angular dependence. This can be implemented using peculiar-velocity catalogues (e.g., CosmicFlows) and, at higher redshift, forthcoming large surveys with wide sky coverage.
More broadly, these results emphasize that cosmological observables are not determined solely by spacetime geometry, but also by the kinematics of the observer congruence used to interpret them. This highlights the importance of a fully covariant, observer-based approach to cosmology, in which the role of the observer is treated as an integral part of the physical description rather than a passive element.
Appendix A Evolution equation of $\sqrt{A}$
Let $A(\lambda)$ be the cross-sectional area of a narrow bundle of null rays, and let the null expansion be defined by
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (A.1)
We define
$X(\lambda)\equiv\sqrt{A(\lambda)},\qquad\Rightarrow\qquad A=X^{2}.$ (A.2)
Differentiating with respect to $\lambda$ gives
$\frac{dA}{d\lambda}=2X\frac{dX}{d\lambda}.$ (A.3)
Substituting (A.3) into (A.1) yields
$\theta_{N}=\frac{1}{X^{2}}\left(2X\frac{dX}{d\lambda}\right)=2\,\frac{1}{X}\frac{dX}{d\lambda}.$ (A.4)
Now differentiate (A.4) once more:
$\frac{d\theta_{N}}{d\lambda}$ $=2\,\frac{d}{d\lambda}\left(\frac{1}{X}\frac{dX}{d\lambda}\right)$
$=2\left[\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}-\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right].$ (A.5)
On the other hand, squaring (A.4) gives
$\frac{1}{2}\,\theta_{N}^{2}=\frac{1}{2}\left(4\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right)=2\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}.$ (A.6)
Adding (A.5) and (A.6), the quadratic terms cancel identically, leaving
$\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=2\,\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}.$ (A.7)
Finally, returning to $X=\sqrt{A}$, we obtain the useful identity
$\boxed{\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=\frac{2}{\sqrt{A}}\frac{d^{2}\sqrt{A}}{d\lambda^{2}}.}$ (A.8)
Appendix B Proof that $d_{A}=aS_{K}(\chi)$ solves the Sachs equation in FLRW
First note that along a radial null geodesic in conformal FLRW coordinates
$ds^{2}=a^{2}(\eta)\left[-d\eta^{2}+d\chi^{2}+S_{K}^{2}(\chi)d\Omega^{2}\right],$
the null condition $ds^{2}=0$ for radial propagation implies
$d\eta^{2}+d\chi^{2}=0\qquad\Rightarrow\qquad\frac{d\chi}{d\eta}=\pm 1.$ (B.1)
For an incoming ray (propagating toward the observer), we take
$\frac{d\chi}{d\eta}=-1.$ (B.2)
By definition of the tangent vector to the null geodesic,
$k^{\eta}=\frac{d\eta}{d\lambda},\qquad k^{\chi}=\frac{d\chi}{d\lambda}.$
Using (B.2), we obtain
$k^{\chi}=\frac{d\chi}{d\eta}\frac{d\eta}{d\lambda}=-\,k^{\eta}.$ (B.3)
Now consider an arbitrary scalar function $f(\eta,\chi)$. The derivative along the null ray is
$\frac{df}{d\lambda}=k^{\alpha}\nabla_{\alpha}f=k^{\eta}\partial_{\eta}f+k^{\chi}\partial_{\chi}f.$ (B.4)
Substituting (B.3), we find
$\frac{df}{d\lambda}=k^{\eta}\partial_{\eta}f-k^{\eta}\partial_{\chi}f=k^{\eta}\left(\partial_{\eta}-\partial_{\chi}\right)f.$ (B.5)
Applying this to $d_{A}=a(\eta)S_{K}(\chi)$ gives
$\frac{dd_{A}}{d\lambda}=k^{\eta}\left(\frac{da}{d\eta}\,S_{K}-a\,\frac{dS_{K}}{d\chi}\right).$ (B.6)
Differentiating once more,
$\frac{d^{2}d_{A}}{d\lambda^{2}}$ $=\frac{d}{d\lambda}\left[k^{\eta}\left(\frac{da}{d\eta}\,S_{K}-a\,\frac{dS_{K}}{d\chi}\right)\right]$
$=\left(\frac{dk^{\eta}}{d\lambda}\right)\left(\frac{da}{d\eta}S_{K}-a\frac{dS_{K}}{d\chi}\right)+(k^{\eta})^{2}\left(\frac{d^{2}a}{d\eta^{2}}S_{K}-2\frac{da}{d\eta}\frac{dS_{K}}{d\chi}+a\frac{d^{2}S_{K}}{d\chi^{2}}\right).$ (B.7)
To eliminate $dk^{\eta}/d\lambda$, we use the $\eta$-component of the null geodesic equation $k^{\beta}\nabla_{\beta}k^{\eta}=0$. In conformal FLRW coordinates the relevant Christoffel symbols are
$\Gamma^{\eta}_{\eta\eta}=\Gamma^{\eta}_{\chi\chi}=\frac{1}{a}\frac{da}{d\eta}.$
Using $(k^{\chi})^{2}=(k^{\eta})^{2}$ for radial null propagation, we obtain
$\frac{dk^{\eta}}{d\lambda}+\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}=0,\qquad\Rightarrow\qquad\frac{dk^{\eta}}{d\lambda}=-\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}.$ (B.8)
Substituting into (B.7), the terms proportional to $\frac{dS_{K}}{d\chi}$ cancel identically, yielding
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}\left[a\frac{d^{2}S_{K}}{d\chi^{2}}+\left(\frac{d^{2}a}{d\eta^{2}}-2\frac{(da/d\eta)^{2}}{a}\right)S_{K}\right].$ (B.9)
Using the identity
$\frac{d^{2}S_{K}}{d\chi^{2}}=-K\,S_{K},$ (B.10)
we obtain
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}S_{K}\left[\frac{d^{2}a}{d\eta^{2}}-\frac{2}{a}\left(\frac{da}{d\eta}\right)^{2}-aK\right]=-(k^{\eta})^{2}\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]d_{A}.$ (B.11)
The Friedmann equations in conformal time read
$\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}+K\right]=\frac{8\pi G}{3}a^{2}\rho+\frac{\Lambda}{3}a^{2},$ (B.12)
and
$\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}=-\frac{4\pi G}{3}a^{2}(\rho+3p)+\frac{\Lambda}{3}a^{2}.$ (B.13)
Combining (B.12)–(B.13) gives the useful identity
$\boxed{\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]=4\pi Ga^{2}(\rho+p).}$ (B.14)
Substituting Eq. (B.14) into Eq. (B.11) we obtain
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-4\pi G\,a^{2}(\rho+p)\,(k^{\eta})^{2}\,d_{A}.$ (B.15)
For a perfect fluid,
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)(u\cdot k)^{2}.$
In conformal FLRW coordinates, comoving observers follow worldlines of constant spatial coordinates. Their four-velocity is aligned with the conformal time direction and must be normalized so that $u^{\alpha}u_{\alpha}=-1$. Since the metric contains an overall factor $a^{2}(\eta)$, this normalization introduces a factor $1/a$, so that the physical four-velocity is proportional to the conformal time direction divided by $a$.
As a result, the photon energy measured by comoving observers is
$(u\cdot k)=a\,k^{\eta}.$
Therefore,
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)\,a^{2}\,(k^{\eta})^{2}.$ (B.16)
Comparing this with Eq. (B.15) immediately gives
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}\,R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A},$ (B.17)
which is precisely the Sachs optical equation.
Appendix C Useful covariant identities
C.1 Homogeneous scalars in FLRW
In an exact FLRW spacetime, the scale factor $a$ depends only on cosmic time and is therefore homogeneous on the spatial hypersurfaces orthogonal to the comoving 4–velocity $u^{\alpha}$. This implies
$h_{\alpha}{}^{\beta}\nabla_{\beta}a=0,$ (C.1)
where $h_{\alpha\beta}=g_{\alpha\beta}+u_{\alpha}u_{\beta}$ is the spatial projector.
Using the standard 1+3 decomposition of the gradient of a scalar,
$\nabla_{\alpha}a=-u_{\alpha}\dot{a}+h_{\alpha}{}^{\beta}\nabla_{\beta}a,\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (C.2)
it follows immediately that
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}.$ (C.3)
C.2 Contractions involving the photon wavevector
The photon wavevector is decomposed relative to the observer congruence as
$k^{\alpha}=E(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,\qquad u_{\alpha}u^{\alpha}=-1.$ (C.4)
From this decomposition one obtains the following useful identities.
First,
$k^{\alpha}k^{\beta}h_{\alpha\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})h_{\alpha\beta}=E^{2}.$ (C.5)
Second, since the shear tensor is spatial, symmetric and orthogonal to $u^{\alpha}$,
$k^{\alpha}k^{\beta}\sigma_{\alpha\beta}=E^{2}\sigma_{\alpha\beta}e^{\alpha}e^{\beta}.$ (C.6)
Third, since the vorticity tensor is antisymmetric,
$k^{\alpha}k^{\beta}\omega_{\alpha\beta}=0.$ (C.7)
Finally, using $u_{\alpha}k^{\alpha}=-E$ and $u_{\alpha}\mathcal{A}^{\alpha}=0$, one finds
$k^{\alpha}k^{\beta}u_{\alpha}\mathcal{A}_{\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})u_{\alpha}\mathcal{A}_{\beta}=-E^{2}\,\mathcal{A}_{\alpha}e^{\alpha}.$ (C.8)
C.3 Relation between affine parameter and observer time
The photon energy measured by an observer with four-velocity $u^{\alpha}$ is defined covariantly as
$E\equiv-(u_{\alpha}k^{\alpha}).$ (C.9)
To relate this quantity to the derivative with respect to the affine parameter, consider the local rest frame of the observer congruence at a given event. In this frame, the four-velocity takes the form
$u^{\alpha}=(1,0,0,0),$ (C.10)
and the photon tangent vector can be written as
$k^{\alpha}=\frac{dx^{\alpha}}{d\lambda}=\left(\frac{dt}{d\lambda},\frac{dx^{i}}{d\lambda}\right).$ (C.11)
Using the definition of $E$, one obtains
$E=-u_{\alpha}k^{\alpha}=\frac{dt}{d\lambda},$ (C.12)
where $t$ denotes the proper time measured by the observer at the event under consideration.
Therefore, locally along the photon trajectory,
$\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}=E\,\frac{d}{dt}.$ (C.13)
This relation is understood as a local statement in the rest frame of the observer congruence. It provides an intuitive interpretation of the affine parameter evolution in terms of the rate at which the photon traverses the observer’s proper time.