### 14 Apr 2026 v2

# Redshift Dipoles from Non-Geodesic Observer Congruences in Covariant Cosmology

Abstract
In an inhomogeneous universe, the identification of the Hubble frame with a geodesic matter flow is not guaranteed beyond the idealized FLRW limit, particularly once structure formation leads to a non-trivial distribution of velocities and gravitational fields. Within a fully covariant framework, we show that a non-geodesic observer congruence introduces an additional contribution to the propagation of redshift along the past light cone, proportional to the line-of-sight projection of the observer four-acceleration. This generates a dipolar modulation in the redshift itself, which propagates to any observable defined in redshift space.
Unlike the standard kinematic dipole associated with a global Lorentz boost, this contribution arises from the kinematics of the observer congruence and depends on its evolution along the light cone, leading to a non-trivial redshift dependence. These results provide a framework to explore potential signatures of non-geodesic kinematics in cosmological data, with possible implications for the interpretation of large-scale dipoles and redshift-based inference.
I Introduction
Cosmological observables are defined on the past light–cone of the observer, reflecting the causal structure of spacetime [5, 7]. Among them, redshift plays a central role, encoding the relation between spacetime dynamics and observational data. In standard cosmology, it is interpreted with respect to the comoving geodesic congruence, for which it reduces to the familiar relation $1+z=a_{0}/a$. However, this interpretation relies on a specific identification between the observer congruence and the underlying matter flow.
In realistic, inhomogeneous spacetimes, the congruence used to define cosmological observables—such as those associated with the Hubble flow or the CMB rest frame—is typically defined through local kinematic corrections and need not correspond to a globally defined geodesic matter flow along the past light cone. This mismatch is expected to become more relevant at low redshift, where structure formation induces significant inhomogeneities, peculiar velocity fields, and anisotropic gravitational environments.
In general, different congruences are related by spacetime-dependent local boosts, under which geodesicity is not preserved. As a result, an observer congruence may possess a non-vanishing four-acceleration even in a spacetime sourced by pressureless matter. Here, the four-acceleration does not describe a force acting on matter, but rather the non-geodesic nature of the observer congruence used to define redshift. This raises a fundamental question: how does the choice of observer congruence, and in particular its possible non-geodesic character, affect the definition and interpretation of cosmological observables? It is important to emphasize that the present analysis does not question the standard practice of applying local kinematic corrections to define observational frames such as the CMB rest frame. Rather, it highlights that such corrections do not uniquely define a geodesic observer congruence along the past light-cone, and therefore need not correspond to a physically global Hubble flow in an inhomogeneous universe. Therefore, the observational relevance of the four-acceleration arises because redshift is operationally defined with respect to this congruence.
In this work, the problem is addressed within a fully covariant framework [5, 16]. The optical focusing equation governing the angular diameter distance is purely geometrical, while the relation between affine parameter and observed redshift depends explicitly on the kinematics of the observer congruence. In particular, a non-geodesic congruence introduces additional contributions to the redshift propagation equation along the past light cone, leading to a dipolar modulation proportional to the projection of the four-acceleration along the line of sight. This result has immediate observational consequences. Since cosmological distances are expressed as functions of redshift, any dipolar modulation in the redshift propagates to all redshift-space observables. Importantly, this dipole originates from the local kinematics of the observer congruence and is not, in general, reducible to a single global Lorentz boost.
An interesting observational point is that accelerated reference frames are not merely a theoretical possibility. The acceleration of the Solar System barycentre produces the well-known secular aberration drift [15, 3], a dipolar proper-motion pattern that has been measured with very long baseline interferometry and more recently constrained with Gaia astrometry. This illustrates that observer acceleration can leave directly observable anisotropic signatures in precision cosmological measurements, and it motivates considering whether analogous kinematic effects may also enter the definition of redshift when the relevant cosmological frame is not geodesic. More broadly, several large-scale dipolar or directional anomalies have been discussed in the literature, including the low-multipole alignments in the CMB and the unexpectedly large dipole reported in quasar number counts (see e.g. [10, 13, 4]). While such observations need not be explained by the mechanism proposed here, they illustrate that the interpretation of cosmic dipoles remains open and deserves careful scrutiny. In that context, the possibility that part of the observed anisotropy may arise from the kinematics of a non-geodesic observer congruence is worth considering, precisely because it follows from the covariant relativistic description itself, without invoking additional degrees of freedom or modifications of gravity.
The structure of this paper is as follows. The Sachs optical equation for the angular diameter distance is first derived, and its purely geometrical character is emphasized. The propagation of redshift in both geodesic and non-geodesic observer congruences is then analyzed, showing that a dipolar modulation naturally arises in the latter case. Order-of-magnitude estimates for the expected amplitude of this effect are presented, and an operational observable designed to isolate the dipolar component in redshift data is introduced. The paper concludes with a discussion of the implications for cosmological inference and possible observational tests.
II The angular diameter distance from the Raychaudhuri equation
Consider a congruence of null geodesics with tangent vector $k^{\alpha}$, affinely parametrised by $\lambda$,
$k^{\beta}\nabla_{\beta}k^{\alpha}=0,\qquad k^{\alpha}k_{\alpha}=0.$ (1)
Let $A(\lambda)$ denote the cross-sectional area of a narrow bundle of rays, defined on the two-dimensional screen space orthogonal to $k^{\alpha}$. The associated null expansion scalar is
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (2)
We define the optical shear and vorticity scalars of the null congruence as
$\sigma_{N}^{2}\equiv\frac{1}{2}\,\sigma^{(N)}_{\alpha\beta}\sigma_{(N)}^{\alpha\beta}\geq 0,\qquad\omega_{N}^{2}\equiv\frac{1}{2}\,\omega^{(N)}_{\alpha\beta}\omega_{(N)}^{\alpha\beta}\geq 0,$ (3)
so that $\sigma^{(N)}_{\alpha\beta}\sigma_{(N)}^{\alpha\beta}=2\sigma_{N}^{2}$ and $\omega^{(N)}_{\alpha\beta}\omega_{(N)}^{\alpha\beta}=2\omega_{N}^{2}$.
The evolution of the null expansion $\theta_{N}$ is governed by the null Raychaudhuri equation [11],
$\frac{d\theta_{N}}{d\lambda}=-\frac{1}{2}\theta_{N}^{2}-2\sigma_{N}^{2}+2\omega_{N}^{2}-R_{\alpha\beta}k^{\alpha}k^{\beta}.$ (4)
With this convention, $\sigma_{N}^{2}$ contributes to focusing, while $\omega_{N}^{2}$ contributes to defocusing.
For a hypersurface-orthogonal null congruence (as appropriate for light bundles forming wavefronts), the vorticity vanishes,
$\omega_{N}^{2}=0.$ (5)
Using $\theta_{N}=(1/A)\,dA/d\lambda$ and the identity (see Appendix A)
$\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\theta_{N}^{2}=\frac{2}{\sqrt{A}}\frac{d^{2}\sqrt{A}}{d\lambda^{2}},$
Raychaudhuri can be rewritten as an evolution equation for $\sqrt{A}$:
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}\left(R_{\alpha\beta}k^{\alpha}k^{\beta}+2\sigma_{N}^{2}\right)\sqrt{A}.$ (6)
For a shear-free light bundle,
$\frac{d^{2}\sqrt{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\sqrt{A}.$ (7)
The angular diameter distance is defined geometrically by
$d_{A}^{2}\equiv\frac{A}{\Omega},$ (8)
where $\Omega$ is the solid angle subtended by the source at the observer. The solid angle is defined in the observer’s local rest frame at the vertex of the light cone and is therefore independent of the affine parameter along the null geodesic. Hence $d_{A}\propto\sqrt{A}$, and the evolution equation for $\sqrt{A}$ in the shear-free case reduces to
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A}.$ (9)
Equation (9) corresponds to the shear-free limit of the Sachs optical equation [12], valid for hypersurface-orthogonal null congruences with vanishing vorticity. In this case, the evolution of the angular diameter distance is governed solely by the Ricci focusing term. The initial conditions at the observer (located at $\lambda=0$) are
$d_{A}(0)=0,\qquad\left.\frac{dd_{A}}{d\lambda}\right|_{0}=1,$ (10)
where the second condition fixes the normalization of the affine parameter such that, locally near the vertex of the light cone, $d_{A}\simeq\lambda$ and $A\simeq\lambda^{2}$. This reflects the locally flat behaviour of the light cone near the observer.
II.1 Matter source term and FLRW universe
Using Einstein’s equations,
$R_{\alpha\beta}=8\pi G\left(T_{\alpha\beta}-\frac{1}{2}Tg_{\alpha\beta}\right)+\Lambda g_{\alpha\beta},$ (11)
and noting that $g_{\alpha\beta}k^{\alpha}k^{\beta}=0$, we obtain
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G\,T_{\alpha\beta}k^{\alpha}k^{\beta}.$ (12)
For a perfect fluid,
$T_{\alpha\beta}=(\rho+p)u_{\alpha}u_{\beta}+pg_{\alpha\beta}.$ (13)
Here $u^{\alpha}$ denotes the four-velocity of the cosmological perfect fluid sourcing the geometry. This implies
$T_{\alpha\beta}k^{\alpha}k^{\beta}=(\rho+p)(u\cdot k)^{2}.$ (14)
Thus the focusing term becomes
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)(u\cdot k)^{2}.$ (15)
Note that $(u\cdot k)$ is the energy $E$ of the photon measured by the observer $u^{\alpha}$ and therefore
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)E^{2}.$ (16)
As a consistency check, specialising to an FLRW spacetime where the null congruence associated with radial light propagation is shear-free and the metric takes its standard form, the solution of Eq. (9) is (see Appendix B)
$d_{A}=a(\eta)\,S_{K}\!\big(\chi\big).$ (17)
with $\eta$ being the conformal time and $\chi$ the comoving distance.
III Redshift definition and the geodesic observer congruence
From this point on, $u^{\alpha}$ will denote the observer congruence used to define measured photon energy and redshift; in exact FLRW it coincides with the fluid 4-velocity. The observed redshift is defined covariantly by
$1+z_{obs}\equiv\frac{(u\cdot k)_{\rm em}}{(u\cdot k)_{\rm obs}},\qquad E\equiv-(u\cdot k),$ (18)
Differentiating $E=-(u\cdot k)$ along the ray gives
$\frac{dE}{d\lambda}=-\,k^{\alpha}\nabla_{\alpha}(u_{\beta}k^{\beta})=-\,k^{\alpha}k^{\beta}\nabla_{\alpha}u_{\beta},$ (19)
where we used $k^{\alpha}\nabla_{\alpha}k^{\beta}=0$ (affine null geodesic). Introducing the standard 1+3 decomposition of $\nabla_{\alpha}u_{\beta}$:
$\nabla_{\alpha}u_{\beta}=\frac{1}{3}\theta\,h_{\alpha\beta}+\sigma_{\alpha\beta}+\omega_{\alpha\beta}-u_{\alpha}\mathcal{A}_{\beta},\qquad h_{\alpha\beta}=g_{\alpha\beta}+u_{\alpha}u_{\beta}.$ (20)
For the photon, decompose $k^{\alpha}$ relative to the observers as
$k^{\alpha}=E\,(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,$ (21)
where $e^{\alpha}$ is the spatial propagation direction measured by $u^{\alpha}$.
Substituting (20) and (21) into (19), using $k^{\alpha}k_{\alpha}=0$ and $h_{\alpha\beta}u^{\beta}=0$, one obtains the exact kinematical propagation law:
$\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta+\sigma_{\alpha\beta}e^{\alpha}e^{\beta}-\mathcal{A}_{\alpha}e^{\alpha}\right).$ (22)
For a geodesic observer congruence,
$\mathcal{A}^{\alpha}\equiv u^{\beta}\nabla_{\beta}u^{\alpha}=0.$ (23)
In the absence of shear in the observer congruence (as in exact FLRW), we have
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,\frac{1}{3}E\theta.}$ (24)
Equation (24) reduces to the standard FLRW redshift law. Indeed, defining the expansion scalar and local Hubble rate of the observer congruence by
$H\equiv\frac{\theta}{3},$ (25)
one has
$\frac{1}{E}\frac{dE}{d\lambda}=-EH.$ (26)
We define a redshift parameter along the null geodesic as $1+z(\lambda)=E(\lambda)/E_{\rm obs}$ along the ray whose value at the source position $\lambda=\lambda_{s}$ coincides with the observed redshift $z_{obs}$ and $E_{\rm obs}$ is constant at the observation event. It follows that
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH.$ (27)
On the other hand, in FLRW the scale factor depends only on cosmic time (see Appendix C), so
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha},\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (28)
Therefore,
$\frac{da}{d\lambda}=k^{\alpha}\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}k^{\alpha}=\dot{a}\,E=aHE.$ (29)
where we used $u_{\alpha}k^{\alpha}=-E$ and $H=\dot{a}/a$. Then
$\frac{1}{a}\frac{da}{d\lambda}=EH.$ (30)
Comparing (27) and (30), we obtain
$\frac{d\ln(1+z)}{d\lambda}=-\frac{d\ln a}{d\lambda},$ (31)
Evaluating at the observer, the last equation integrates to
$1+z_{obs}=\frac{a_{0}}{a},$ (32)
where $a$ is the scale factor evaluated at the source position. Thus, Eq. (22) provides the exact covariant propagation law whose geodesic, shear-free FLRW limit reproduces the standard cosmological redshift relation.
IV Non-geodesic observer congruence and the dipole modulation
The Sachs equation (9) for $d_{A}$ is purely geometrical and does not require the observer congruence to be geodesic. In the FLRW derivation above, the assumption of geodesicity entered only when relating the affine parameter (or conformal time) to the redshift parameter $z$ defined along the null geodesic. We now relax this assumption and make explicit where and how the modification appears.
In the previous FLRW derivation we implicitly took $u^{\alpha}$ to be the comoving geodesic congruence. We now relax (23) and allow
$\mathcal{A}^{\alpha}\neq 0.$ (33)
The impact of (33) is entirely contained in the propagation of $E$ along the null ray. In full generality, the propagation law is
$\boxed{\frac{1}{E}\frac{dE}{d\lambda}=-\,E\left(\frac{1}{3}\theta+\sigma_{\alpha\beta}e^{\alpha}e^{\beta}-\mathcal{A}_{\alpha}e^{\alpha}\right).}$ (34)
The different kinematical contributions have a clear angular structure: the expansion scalar $\theta$ contributes to the monopole, the projection $\mathcal{A}_{\alpha}e^{\alpha}$ generates a dipolar term, and the shear contribution $\sigma_{\alpha\beta}e^{\alpha}e^{\beta}$ induces a quadrupolar modulation.
Equation (34) can be written as
$\frac{1}{E}\frac{dE}{d\lambda}=-EH-E\,\sigma_{\alpha\beta}e^{\alpha}e^{\beta}+E\,\mathcal{A}_{\alpha}e^{\alpha},$ (35)
where $H=\theta/3$. Using $1+z=E/E_{\rm obs}$, with $E_{\rm obs}$ constant, we obtain
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH-E\,\sigma_{\alpha\beta}e^{\alpha}e^{\beta}+E\,\mathcal{A}_{\alpha}e^{\alpha}.$ (36)
In the present work, we focus on the direction-dependent modulation of the redshift and therefore restrict attention to the anisotropic contributions. Under this assumption, and keeping the FLRW background relation for the scale factor evolution (30), the standard relation $(1+z)a=a_{0}$ is replaced by,
$(1+z_{\rm obs})=\frac{a_{0}}{a}\exp\left[\int_{\lambda_{0}}^{\lambda_{s}}E\left(\mathcal{A}_{\alpha}e^{\alpha}-\sigma_{\alpha\beta}e^{\alpha}e^{\beta}\right)d\lambda\right].$ (37)
For weak departures from the geodesic FLRW case, this becomes
$1+z_{\rm obs}\simeq\frac{a_{0}}{a}\left[1+\int_{\lambda_{0}}^{\lambda_{s}}E\left(\mathcal{A}_{\alpha}e^{\alpha}-\sigma_{\alpha\beta}e^{\alpha}e^{\beta}\right)d\lambda\right].$ (38)
Since $\mathcal{A}_{\alpha}e^{\alpha}$ is the projection of a spatial vector along the line of sight, it has the angular structure of a dipole. Likewise, the contraction $\sigma_{\alpha\beta}e^{\alpha}e^{\beta}$ is the projection of a symmetric trace-free tensor and therefore has the angular structure of a quadrupole. This shows that non-geodesicity of the observer congruence naturally induces a dipolar modulation in the redshift itself, while anisotropic expansion encoded in the shear contributes a quadrupolar component. Therefore, the dipole and quadrupole anisotropies in redshift-space observables may provide complementary probes of the acceleration and shear of the observer congruence, respectively.
The non-geodesic dipole arises from a line-of-sight integral of the observer acceleration field and therefore depends on the photon trajectory and redshift. This contrasts with the standard kinematic dipole, which is generated by a constant boost and is independent of distance. As a result, this dipole cannot, in general, be removed by a single global Lorentz transformation, since it arises from an integrated, position-dependent acceleration field rather than from a constant boost.
The shear term represents the anisotropic expansion of the chosen observer congruence. In an exact FLRW description, this term vanishes identically. However, in realistic inhomogeneous settings, an observer congruence used to interpret cosmological observables need not remain shear-free along the past light cone. In this sense, the quadrupolar contribution can be interpreted as encoding deviations from an idealized isotropic Hubble flow, potentially associated with the anisotropic kinematics of large-scale structure. This suggests a possible connection between low-redshift quadrupolar anisotropies in distance-redshift observables [14] and the anisotropic kinematics of nearby large-scale structure [6]. Exploring this observational connection in detail is beyond the scope of the present work and is left for future study.
V Upper Bound Estimate of the Non–Geodesic Dipole
The dipolar contribution associated with a non–geodesic observer congruence arises from the term $\mathcal{A}_{\alpha}e^{\alpha}$ in the covariant propagation equation 34. As discussed in the previous section, the different kinematical terms have distinct angular structures: the expansion scalar contributes to the monopole, the acceleration term $\mathcal{A}_{\alpha}e^{\alpha}$ generates a dipole, and the shear term $\sigma_{\alpha\beta}e^{\alpha}e^{\beta}$ induces a quadrupole. In this section, we focus on isolating the dipolar contribution associated with $\mathcal{A}_{\parallel}\equiv\mathcal{A}_{\alpha}e^{\alpha}$,
$\frac{d}{d\lambda}\ln(1+z)=-E\left(\frac{1}{3}\theta-\mathcal{A}_{\alpha}e^{\alpha}\right).$ (39)
To estimate the amplitude of the anisotropic modulation, we separate the monopole contribution and define
$\ln(1+z)\simeq\ln(1+\bar{z})+\delta\ln(1+z)_{\rm ng},$ (40)
where $\bar{z}$ is the monopolar redshift, satisfying
$\frac{d}{d\lambda}\ln(1+\bar{z})=-EH.$ (41)
The dipolar residual is then governed by
$\left.\frac{d}{d\lambda}\,\delta\ln(1+z)\right|_{\rm ng}\simeq E\,\mathcal{A}_{\parallel}.$ (42)
To express this in a more intuitive form, recall that along the photon trajectory
$k^{\alpha}=\frac{dx^{\alpha}}{d\lambda},$ (43)
and therefore
$E=-u_{\alpha}k^{\alpha}=\frac{dt}{d\lambda}.$ (44)
Using
$\frac{d}{d\lambda}=E\frac{d}{dt},$ (45)
Eq. (42) becomes
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\simeq\mathcal{A}_{\parallel},$ (46)
so that, restoring dimensions,
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{c}.$ (47)
Integrating along the photon path gives
$\delta\ln(1+z_{obs})_{\rm ng}\sim\int\frac{\mathcal{A}_{\parallel}}{c}\,dt.$ (48)
At this stage, it is important to emphasize that the evolution of the effective four-acceleration associated with an observer congruence in an inhomogeneous universe is currently a subject of ongoing debate. Different approaches, particularly relativistic and quasi–Newtonian treatments of large–scale structure formation, have led to qualitatively different conclusions regarding both the magnitude and time dependence of $\mathcal{A}_{a}$ (see e.g. [8, 2, 18]). In this work, we do not attempt to model the detailed evolution of $\mathcal{A}_{\parallel}(t)$. Instead, we adopt a conservative phenomenological approach aimed at estimating the possible amplitude of the induced dipole.
As a simplifying approximation, we consider the case where $\mathcal{A}_{\parallel}$ can be treated as approximately constant over a restricted low–redshift interval. This assumption over a restricted interval should be interpreted as an effective bound on the line-of-sight integral, rather than as a statement about the time evolution of the acceleration itself.
Under this approximation,
$\delta\ln(1+z_{obs})_{\rm ng}\lesssim\frac{\mathcal{A}_{\parallel}}{c}\,\Delta t\approx\frac{\mathcal{A}_{\parallel}L}{c^{2}},$ (49)
where $L\simeq c\,\Delta t$ is the effective light–travel distance.
At low redshift, the light-travel distance is of order the Hubble-law distance,
$L\simeq\frac{c}{H_{0}}\,z.$ (50)
Substituting Eq. (50) into Eq. (49) yields
$\delta\ln(1+z_{obs})_{\rm ng}\lesssim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (51)
For $z\ll 1$, one has $\delta\ln(1+z)\simeq\delta z$, so that
$\delta z_{\rm ng}\lesssim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (52)
This relation can be recast as
$\mathcal{A}_{\parallel}\gtrsim\frac{cH_{0}}{z}\,\delta z_{\rm ng}.$ (53)
Taking $H_{0}\simeq 70\,{\rm km\,s^{-1}\,Mpc^{-1}}$, one finds
$cH_{0}\simeq 6.8\times 10^{-10}\ {\rm m\,s^{-2}}.$ (54)
At $z\simeq 0.1$, representative of the shallow redshift regime relevant for peculiar-velocity surveys, this gives
$\mathcal{A}_{\parallel}\gtrsim 6.8\times 10^{-9}\,\delta z_{\rm ng}\quad{\rm m\,s^{-2}}.$ (55)
Therefore, dipole amplitudes in the range $\delta z_{\rm ng}\sim 10^{-3}$$10^{-5}$ correspond to accelerations
$\mathcal{A}_{\parallel}\gtrsim 10^{-12}-10^{-14}\ {\rm m\,s^{-2}}.$ (56)
For comparison, one may consider characteristic rates of change of peculiar velocities associated with large-scale structure formation, as discussed in both Newtonian and relativistic treatments (see e.g. [9]). These analyses suggest that the evolution of peculiar velocities occurs over cosmological timescales, leading to a dimensional scaling of the form
$\mathcal{A}_{\rm LSS}\sim vH_{0},$ (57)
where $v\sim 300\,{\rm km\,s^{-1}}$ is a typical peculiar velocity amplitude.
This scaling should be understood as a dimensional proxy for the rate of change of velocities over cosmological timescales, rather than as a prediction of a coherent acceleration field or as implying $\mathcal{A}_{\rm LSS}\sim vH_{0}$ at all times. In realistic structure formation scenarios, the evolution of peculiar velocities is driven by local gravitational potentials and is neither spatially uniform nor constant in time. This gives
$\mathcal{A}_{\rm LSS}\sim 7\times 10^{-13}\ {\rm m\,s^{-2}}.$ (58)
Using this as a reference scale, one finds that dipolar modulations of order $\delta z\lesssim 10^{-4}$ at $z\sim 0.1$ are compatible with this dimensional estimate, although a realistic prediction requires specifying the temporal evolution and coherence properties of the effective acceleration field.
The estimates presented here should be interpreted as order-of-magnitude upper bounds over a restricted low-redshift interval, aimed at assessing the potential significance of the effect rather than providing a detailed physical model.
VI Observational Implementation and Testable Signatures
In a perturbed FLRW spacetime, peculiar velocities of sources and observers produce a Doppler dipole. If a source moves with a small peculiar velocity $v_{s}^{\alpha}$ relative to the cosmological frame $u^{\alpha}$, and the observer has velocity $v_{o}^{\alpha}$, the observed redshift at linear order in peculiar velocities can be written as
$1+z\simeq(1+\bar{z})\left[1+(v_{s}-v_{o})_{\alpha}e^{\alpha}\right],$ (59)
where $\bar{z}$ denotes the purely monopolar redshift generated by the background expansion.
Combining this with the covariant propagation law derived above and for weak non-geodesic effects, the dipolar component of the redshift field can schematically be written as
$\delta z\sim(1+\bar{z})\left[(v_{s}-v_{o})_{\alpha}e^{\alpha}+\int_{\lambda_{o}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right],$ (60)
where the first term represents the standard kinematic contribution from peculiar velocities, while the second term arises from the non–geodesic nature of the observer congruence.
Both contributions share the same angular structure (dipolar), but differ in their physical origin and redshift dependence. The kinematic term is local and reflects peculiar velocities at emission and observation, whereas the non–geodesic contribution is an integrated effect along the photon trajectory. This distinction provides the basis for an observational test. In peculiar velocity surveys such as CosmicFlows [19], redshifts are typically corrected to the CMB rest frame, removing the dominant contribution from the observer’s local motion. In this case, the remaining dipole is expected to be dominated by large–scale structure (LSS) kinematics, with a possible additional contribution arising from non–geodesic effects.
At low redshift, this motivates a phenomenological model for the dipolar modulation of the redshift field of the form
$\delta z(\bar{z})=\,\hat{n}\cdot\mathbf{D}(\bar{z}),$ (61)
where $\bar{z}$ is the isotropic (monopole) component and $\mathbf{D}(\bar{z})=D(\bar{z})\,\hat{p}(\bar{z})$ is the dipole vector, with $D(\bar{z})$ its amplitude and $\hat{p}(\bar{z})$ its direction.
In this framework, the standard kinematic dipole is expected to reflect the coherence scale and amplitude of peculiar velocities, and therefore to exhibit a redshift dependence determined by large–scale structure dynamics. In contrast, a non–geodesic contribution introduces an additional component sourced by a line–of–sight integral of the observer acceleration field, which need not follow the same scaling. As a consequence, the two contributions need not follow the same redshift evolution. Phenomenologically, one may write
$\mathbf{D}(\bar{z})=\mathbf{D}_{\rm kin}(\bar{z})+\mathbf{D}_{\rm ng}(\bar{z}),$ (62)
where the two contributions need not be aligned and need not follow the same redshift evolution.
Therefore, by measuring the redshift dependence of the dipole amplitude and its directional stability, one can test whether the observed signal is consistent with purely kinematic expectations or requires an additional non–geodesic contribution. In practice, detecting this effect requires disentangling it from observational uncertainties and intrinsic scatter in peculiar velocity measurements, which may dominate at low redshift. Overall, this provides a concrete observational avenue to test whether the redshift dipole is fully accounted for by large–scale structure kinematics, or whether additional contributions are present.
VII Discussion and Conclusions
In this work, we have shown that non-geodesic observer congruences generically induce a dipolar modulation in the cosmological redshift through the projection of the four-acceleration along the line of sight. This effect arises directly at the level of the redshift and therefore propagates to any observable expressed in redshift space. Since redshift is operationally defined with respect to a chosen observational frame (e.g. the CMB frame), this contribution corresponds to a physically distinct inference, rather than to a coordinate artifact or a trivial reparametrization.
A key aspect of this result is that the induced dipole is not equivalent to the standard kinematic dipole generated by a constant peculiar velocity. In the latter case, the dipole can be removed by a single global Lorentz boost. In contrast, the contribution derived here is sourced by the generally position-dependent acceleration of the observer congruence and accumulates along the past lightcone. As a result, it cannot, in general, be removed by a global boost transformation and instead represents a genuinely non-local contribution to the observed redshift.
It is important to emphasize that this does not challenge the standard practice of applying local kinematic corrections to define observational frames such as the CMB rest frame. Rather, it highlights that such corrections do not uniquely define a geodesic observer congruence along the past lightcone, and therefore need not correspond to a physically global Hubble flow in an inhomogeneous universe. This highlights that redshift is not a purely geometric observable, but a kinematically defined quantity that depends on the choice of observer congruence. Consequently, cosmological inference based on redshift measurements implicitly inherits assumptions about the underlying kinematics of the chosen observational frame.
The covariant formulation also shows that different kinematical quantities generate distinct angular structures. While the acceleration term produces a dipolar modulation, the shear contribution induces a quadrupolar component. The latter could be related with anisotropic large-scale structure and may be related to low-redshift quadrupolar patterns in distance-redshift observables. A quantitative observational study of this connection is beyond the scope of the present work and will be explored in future analyses.
These considerations are particularly relevant in the low-redshift regime, where structure formation leads to significant inhomogeneities, peculiar velocity fields, and anisotropic environments. In this context, any effective non-geodesic contribution should be interpreted as a local and evolving property of the observer congruence. Accordingly, the estimates presented in this work are best regarded as order-of-magnitude upper bounds over a restricted low-redshift interval, rather than as a description of a coherent acceleration field or a detailed dynamical model.
The distinction presented here is closely related, but not equivalent, to previous work on tilted cosmologies, where observer congruences moving with respect to the matter flow exhibit modified kinematical properties (see e.g. [17, 1]). In those approaches, the focus is typically on local frame transformations and their impact on dynamical variables. By contrast, the effect derived here arises from the propagation of redshift along the past lightcone and therefore encodes an integrated contribution of the observer kinematics that is not captured by local boosts alone.
From an observational perspective, this framework suggests that the relevant signature is not only the amplitude of a dipolar component, but also its redshift evolution and directional behaviour. A purely kinematic dipole associated with peculiar velocities is expected to trace the structure of the local velocity field, whereas a non-geodesic contribution may exhibit a distinct redshift dependence or a misalignment with respect to the reconstructed kinematic dipole. These features could, in principle, be tested using low-redshift peculiar-velocity catalogues (e.g. CosmicFlows) as well as distance-redshift data.
The upper bound estimates indicate that the acceleration required to produce observable dipolar effects is comparable, at the level of dimensional analysis, to characteristic acceleration scales associated with large-scale structure formation. This suggests that the effect may be relevant in low-redshift datasets, where peculiar velocities and bulk flows are significant, although a quantitative prediction requires specifying the temporal evolution and coherence properties of the effective acceleration field. A quantitative assessment of these effects requires extending the present framework to realistic cosmological models and datasets, including forward modeling of redshift-space observables and their angular structure. This could be implemented using peculiar-velocity catalogues and, at higher redshift, forthcoming wide-field surveys.
It is interesting to note that the effect derived here is purely kinematical, yet it raises the question of how it relates to more familiar integrated contributions such as gravitational redshift effects in inhomogeneous spacetimes. While the present formulation does not rely on a potential-based description, clarifying the connection between these viewpoints could provide further insight into the physical origin of redshift anisotropies and deserves separate investigation.
More broadly, these results highlight that cosmological observables are determined not only by spacetime geometry, but also by the kinematics of the observer congruence used to interpret them. This reinforces the importance of a fully covariant, observer-based approach to cosmology, in which the role of the observer is treated as an integral part of the physical description.
Appendix A Evolution equation of $\sqrt{A}$
Let $A(\lambda)$ be the cross-sectional area of a narrow bundle of null rays, and let the null expansion be defined by
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (63)
We define
$X(\lambda)\equiv\sqrt{A(\lambda)},\qquad\Rightarrow\qquad A=X^{2}.$ (64)
Differentiating with respect to $\lambda$ gives
$\frac{dA}{d\lambda}=2X\frac{dX}{d\lambda}.$ (65)
Substituting (65) into (63) yields
$\theta_{N}=\frac{1}{X^{2}}\left(2X\frac{dX}{d\lambda}\right)=2\,\frac{1}{X}\frac{dX}{d\lambda}.$ (66)
Now differentiate (66) once more:
$\frac{d\theta_{N}}{d\lambda}$ $=2\,\frac{d}{d\lambda}\left(\frac{1}{X}\frac{dX}{d\lambda}\right)$
$=2\left[\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}-\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right].$ (67)
On the other hand, squaring (66) gives
$\frac{1}{2}\,\theta_{N}^{2}=\frac{1}{2}\left(4\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right)=2\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}.$ (68)
Adding (67) and (68), the quadratic terms cancel identically, leaving
$\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=2\,\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}.$ (69)
Finally, returning to $X=\sqrt{A}$, we obtain the useful identity
$\boxed{\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=\frac{2}{\sqrt{A}}\frac{d^{2}\sqrt{A}}{d\lambda^{2}}.}$ (70)
Appendix B Proof that $d_{A}=aS_{K}(\chi)$ solves the Sachs equation in FLRW
First note that along a radial null geodesic in conformal FLRW coordinates
$ds^{2}=a^{2}(\eta)\left[-d\eta^{2}+d\chi^{2}+S_{K}^{2}(\chi)d\Omega^{2}\right],$
the null condition $ds^{2}=0$ for radial propagation implies
$d\eta^{2}+d\chi^{2}=0\qquad\Rightarrow\qquad\frac{d\chi}{d\eta}=\pm 1.$ (71)
For an incoming ray (propagating toward the observer), we take
$\frac{d\chi}{d\eta}=-1.$ (72)
By definition of the tangent vector to the null geodesic,
$k^{\eta}=\frac{d\eta}{d\lambda},\qquad k^{\chi}=\frac{d\chi}{d\lambda}.$
Using (72), we obtain
$k^{\chi}=\frac{d\chi}{d\eta}\frac{d\eta}{d\lambda}=-\,k^{\eta}.$ (73)
Now consider an arbitrary scalar function $f(\eta,\chi)$. The derivative along the null ray is
$\frac{df}{d\lambda}=k^{\alpha}\nabla_{\alpha}f=k^{\eta}\partial_{\eta}f+k^{\chi}\partial_{\chi}f.$ (74)
Substituting (73), we find
$\frac{df}{d\lambda}=k^{\eta}\partial_{\eta}f-k^{\eta}\partial_{\chi}f=k^{\eta}\left(\partial_{\eta}-\partial_{\chi}\right)f.$ (75)
Applying this to $d_{A}=a(\eta)S_{K}(\chi)$ gives
$\frac{dd_{A}}{d\lambda}=k^{\eta}\left(\frac{da}{d\eta}\,S_{K}-a\,\frac{dS_{K}}{d\chi}\right).$ (76)
Differentiating once more,
$\frac{d^{2}d_{A}}{d\lambda^{2}}$ $=\frac{d}{d\lambda}\left[k^{\eta}\left(\frac{da}{d\eta}\,S_{K}-a\,\frac{dS_{K}}{d\chi}\right)\right]$
$=\left(\frac{dk^{\eta}}{d\lambda}\right)\left(\frac{da}{d\eta}S_{K}-a\frac{dS_{K}}{d\chi}\right)+(k^{\eta})^{2}\left(\frac{d^{2}a}{d\eta^{2}}S_{K}-2\frac{da}{d\eta}\frac{dS_{K}}{d\chi}+a\frac{d^{2}S_{K}}{d\chi^{2}}\right).$ (77)
To eliminate $dk^{\eta}/d\lambda$, we use the $\eta$-component of the null geodesic equation $k^{\beta}\nabla_{\beta}k^{\eta}=0$. In conformal FLRW coordinates the relevant Christoffel symbols are
$\Gamma^{\eta}_{\eta\eta}=\Gamma^{\eta}_{\chi\chi}=\frac{1}{a}\frac{da}{d\eta}.$
Using $(k^{\chi})^{2}=(k^{\eta})^{2}$ for radial null propagation, we obtain
$\frac{dk^{\eta}}{d\lambda}+\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}=0,\qquad\Rightarrow\qquad\frac{dk^{\eta}}{d\lambda}=-\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}.$ (78)
Substituting into (77), the terms proportional to $\frac{dS_{K}}{d\chi}$ cancel identically, yielding
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}\left[a\frac{d^{2}S_{K}}{d\chi^{2}}+\left(\frac{d^{2}a}{d\eta^{2}}-2\frac{(da/d\eta)^{2}}{a}\right)S_{K}\right].$ (79)
Using the identity
$\frac{d^{2}S_{K}}{d\chi^{2}}=-K\,S_{K},$ (80)
we obtain
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}S_{K}\left[\frac{d^{2}a}{d\eta^{2}}-\frac{2}{a}\left(\frac{da}{d\eta}\right)^{2}-aK\right]=-(k^{\eta})^{2}\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]d_{A}.$ (81)
The Friedmann equations in conformal time read
$\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}+K\right]=\frac{8\pi G}{3}a^{2}\rho+\frac{\Lambda}{3}a^{2},$ (82)
and
$\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}=-\frac{4\pi G}{3}a^{2}(\rho+3p)+\frac{\Lambda}{3}a^{2}.$ (83)
Combining (82)–(83) gives the useful identity
$\boxed{\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]=4\pi Ga^{2}(\rho+p).}$ (84)
Substituting Eq. (84) into Eq. (81) we obtain
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-4\pi G\,a^{2}(\rho+p)\,(k^{\eta})^{2}\,d_{A}.$ (85)
For a perfect fluid,
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)(u\cdot k)^{2}.$
In conformal FLRW coordinates, comoving observers follow worldlines of constant spatial coordinates. Their four-velocity is aligned with the conformal time direction and must be normalized so that $u^{\alpha}u_{\alpha}=-1$. Since the metric contains an overall factor $a^{2}(\eta)$, this normalization introduces a factor $1/a$, so that the physical four-velocity is proportional to the conformal time direction divided by $a$.
As a result, the photon energy measured by comoving observers is
$(u\cdot k)=a\,k^{\eta}.$
Therefore,
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)\,a^{2}\,(k^{\eta})^{2}.$ (86)
Comparing this with Eq. (85) immediately gives
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}\,R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A},$ (87)
which is precisely the Sachs optical equation.
Appendix C Useful covariant identities
C.1 Homogeneous scalars in FLRW
In an exact FLRW spacetime, the scale factor $a$ depends only on cosmic time and is therefore homogeneous on the spatial hypersurfaces orthogonal to the comoving 4–velocity $u^{\alpha}$. This implies
$h_{\alpha}{}^{\beta}\nabla_{\beta}a=0,$ (88)
where $h_{\alpha\beta}=g_{\alpha\beta}+u_{\alpha}u_{\beta}$ is the spatial projector.
Using the standard 1+3 decomposition of the gradient of a scalar,
$\nabla_{\alpha}a=-u_{\alpha}\dot{a}+h_{\alpha}{}^{\beta}\nabla_{\beta}a,\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (89)
it follows immediately that
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}.$ (90)
C.2 Contractions involving the photon wavevector
The photon wavevector is decomposed relative to the observer congruence as
$k^{\alpha}=E(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,\qquad u_{\alpha}u^{\alpha}=-1.$ (91)
From this decomposition one obtains the following useful identities.
First,
$k^{\alpha}k^{\beta}h_{\alpha\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})h_{\alpha\beta}=E^{2}.$ (92)
Second, since the shear tensor is spatial, symmetric and orthogonal to $u^{\alpha}$,
$k^{\alpha}k^{\beta}\sigma_{\alpha\beta}=E^{2}\sigma_{\alpha\beta}e^{\alpha}e^{\beta}.$ (93)
Third, since the vorticity tensor is antisymmetric,
$k^{\alpha}k^{\beta}\omega_{\alpha\beta}=0.$ (94)
Finally, using $u_{\alpha}k^{\alpha}=-E$ and $u_{\alpha}\mathcal{A}^{\alpha}=0$, one finds
$k^{\alpha}k^{\beta}u_{\alpha}\mathcal{A}_{\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})u_{\alpha}\mathcal{A}_{\beta}=-E^{2}\,\mathcal{A}_{\alpha}e^{\alpha}.$ (95)
C.3 Relation between affine parameter and observer time
The photon energy measured by an observer with four-velocity $u^{\alpha}$ is defined covariantly as
$E\equiv-(u_{\alpha}k^{\alpha}).$ (96)
To relate this quantity to the derivative with respect to the affine parameter, consider the local rest frame of the observer congruence at a given event. In this frame, the four-velocity takes the form
$u^{\alpha}=(1,0,0,0),$ (97)
and the photon tangent vector can be written as
$k^{\alpha}=\frac{dx^{\alpha}}{d\lambda}=\left(\frac{dt}{d\lambda},\frac{dx^{i}}{d\lambda}\right).$ (98)
Using the definition of $E$, one obtains
$E=-u_{\alpha}k^{\alpha}=\frac{dt}{d\lambda},$ (99)
where $t$ denotes the proper time measured by the observer at the event under consideration.
Therefore, locally along the photon trajectory,
$\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}=E\,\frac{d}{dt}.$ (100)
This relation is understood as a local statement in the rest frame of the observer congruence. It provides an intuitive interpretation of the affine parameter evolution in terms of the rate at which the photon traverses the observer’s proper time.